# Invar-RAG: Invariant LLM-aligned Retrieval for Better Generation

### v1

@@ 3 Experiment
3.1 Setting
3.1.1 Datasets
Following the prior works Lin et al. (2023b); Asai et al. (2023), we choose two ODQA datasets (FreebaseQA Yao et al. (2014) and MS-MARCO Bajaj et al. (2016)) and one reading comprehension (RC) dataset to do the representation learning in the retrieval stage (denoted as $\mathcal{D}_{R}$) while leveraging other three datasets (Web Question Dumais et al. (2002), Wiki Question Answering Yang et al. (2015)) and SQuAD v2111https://rajpurkar.github.io/SQuAD-explorer/explore/v2.0/dev/Prime_number.html for fine-tuning the LLM in the generation stage (denoted as $\mathcal{D}_{G}$). The statistic of chosen datasets is shown in Tab.1. For detailed descriptions and complied templates, please refer to Appendix A.
Table 1: The Statistics of Fine-tuning Datasets.
@@ 3.1.1 Datasets
Wiki QA Yang et al. (2015) wiki_qa 20360 Open-domain QA
FreebaseQA Yao et al. (2014) freebase_qa 20358 Open-domain QA
MS-MARCO Bajaj et al. (2016) ms_marco 80143 Open-domain QA
Web Question Dumais et al. (2002) web_question 3778 Open-domain QA
@@
3.1.2 Evaluation
To access our performance, we conduct the evaluation on four knowledge-intensive datasets, such as i.e., TriviaQA (denoted as TQA) [^2] https://nlp.cs.washington.edu/triviaqa, Natural Question (denoted as NQ) [^3] https://ai.google.com/research/NaturalQuestions and PopQA [^4] https://huggingface.co/datasets/akariasai/PopQA, that are not involved in the training progress. For the evaluation metric, we evaluate our model’s generation performance using the Exact Match Wongsuphasawat et al. (2012), which indicates whether gold answers are included in the model generations followed by the setting in prior work Lin et al. (2023b); Mao et al. (2024). Furthermore, to evaluate our proposed retriever’s performance, we employ the Acc@5 and Acc@20 as evaluation metrics, which are widely used in related studies Chen et al. (2024); Izacard et al. (2021). These metrics assess the proportion of questions where the correct answers appear in the top-5 or top-20 retrieval results, offering a comprehensive evaluation of the retrieval performance. For more details, please refer to the description and methods in Appdendix B.
3.1.3 Implementation Details
@@
Retrieval Stage. Following the previous work’s setting Ma et al. (2024), as LLaMA is a decoder-only architecture, we append an end-of-sequence token <EOS> to the input sequence and regard the last layer representation as the dense representation to calculate the similarity score. Considering the possible effect caused by the size of each dense representation, we also employ the normalization procedure to map the original representation into unit vectors during both the training and inference stages. For the fine-tuning progress in the retrieval stage, we adopt LoRA architecture Hu et al. (2021) to reduce the high cost of GPU memory. The detailed hyperparameters we used can be found in Appendix C.
Generation Stage. We hold $\overline{k}$ in Sec.2.3 equal to 5 to generate instances for a single example and append multiple examples together to improve the efficiency (the length is limited to 4096 tokens). The used hyperparameters are also shown in the Appendix C. Other implementation details are the same as original papers Lin et al. (2023b); Shi et al. (2023).
3.1.4 Baselines
To demonstrate the effectiveness of our proposed architecture, we compare the retrieval performance of our Invar-RAG with state-of-the-art retrieval methods, including sparse retrieval (BM25 Ram et al. (2023)), dense retrieval (BGE Xiao et al. (2024), Contriever Izacard et al. (2021)) and LLM-based retrieval (LLM-embedder Zhang et al. (2023) and RepLLaMA Ma et al. (2024)). Furthermore, for the corresponding RAG performance, we conduct extensive experiments compared to the novel retriever + generation model to show our superiority. The descriptions for each baseline are listed in the Appendix D.
3.2 Overall Performance
@@
Table 2: Retrieval performance comparison between our designed retriever in Invar-RAG and other baselines. The best results are bold, and the second-best are underlined.
@@
BM25 Ram et al. (2023) 62.5 73.0 49.0 67.0 35.5 51.5
BM25+BGE(re-ranker) Chen et al. (2024) 72.5 78.0 68.0 76.5 54.0 60.0
Contriever Izacard et al. (2021) 68.0 80.5 68.0 84.0 62.0 77.5
BGE-base Xiao et al. (2024) 69.5 80.0 77.0 86.0 72.0 83.0
LLM-embedder Zhang et al. (2023) 67.5 77.5 75.5 86.5 70.0 79.5
RepLLaMA Ma et al. (2024) 66.5 76.0 72.0 85.5 68.5 74.5
@@
3.2.1 Retrieval Performance
In this section, we will present and analyze the retrieval performance of our designed architecture. As illustrated in Tab.2, the sparse retriever BM25 fails to map the given text to proper representations. Although employing an additional model as the re-ranker improves the performance to some extent, the retrieval capability remains sub-optimal due to the inferiority of BM25. Besides, Novel dense retrievers, like BGE and Contriever, present comparable performance over the three datasets, suggesting their effectiveness in leveraging contrastive learning or task-specific fine-tuning. However, they still slightly lag behind our designed Invar-retrieval because of the neglect of rich semantic information Ma et al. (2024). Current researchers have proposed several LLM-based retrievers i.e., LLM-embedder Zhang et al. (2023) and RepLLaMA Ma et al. (2024), which leverage the rich prior knowledge that LLM initially has. However, due to the high cost of processing the massive corpus, it is infeasible to handle all the chunks within the LLM. Moreover, the variance problem that happens in LLM also leads to relatively inferior performance. Correspondingly, we propose our LLM-based retrieval model, Invar-retrieval, as a part of our designed Invar-RAG. The results shows that our methods outperform all the sparse, dense and LLM-based retrievers, especially under the Acc@5 measurement, contributing to our designed invariance loss in reducing the variant and ineffective patterns.
3.2.2 Generation Performance
@@
Table 3: Generation performance comparison between our designed Invar-RAG and other baselines. The best results are bold, and the second-best are underlined.
@@ 3.3 Ablation Study
Table 4: The Statistics of Datasets in Ablation study.
@@
Table 5: Ablation Study on TQA.
@@ 3.4 Invariance Analysis
  • Prior works Zhang et al. (2024) have shown that adding a suitable amount of irrelevant or relatively ineffective documents does help improve the retrieval performance.
Figure 2: Special example for illustrating the effectiveness of invariance loss.
@@
4 Related Work
Information Retrieval: Advancements in deep learning have revolutionized information retrieval systems, enhancing their personalization and accuracy in retrieving relevant documents. Early information retrieval frameworks employed sparse retrievers Ram et al. (2023) or dense retrievers Izacard et al. (2021); Xiao et al. (2024) to represent large corpora but struggled to capture deep semantic relationships Karpukhin et al. (2020). LLM-based retrievers (generative retrieval) have since emerged as notable methods, leveraging the rich prior knowledge of LLMs to significantly improve performance by converting documents into parametric knowledge and generating them instead of computing similarity scores Zhu et al. (2021). However, the frequent encoding and decoding processes in LLMs severely hinder efficiency Zhu et al. (2021). To address the trade-off between effectiveness and efficiency, we propose invar-retrieval in our architecture, enabling the model to efficiently retrieve the most relevant documents without introducing variance.
Retrieval-augmented Language Model: Currently, retrieval-augmented language models have proven effective in answering questions by leveraging external information through the integration of novel retrievers and LLMs Zhu et al. (2021). However, the architectural gap between retrieval and generation continues to hinder unified optimization across the entire retrieval-augmented generation system Fan et al. (2024). To address the isolation between retrieval and generation, a novel architecture called RA-DIT was introduced Lin et al. (2023b). By aligning retriever scoring with LSR scoring Shi et al. (2023), it has been shown to deliver state-of-the-art performance across various tasks. However, it still employs dense retrievers like DRAGON+ Lin et al. (2023a) in the retrieval stage, which fails to eliminate the problem at its source and introduces inefficiencies throughout the process. Correspondingly, we introduce a representation learning method and invariance loss in our Invar-RAG architecture, which partially addresses these issues and explores a novel approach to using a single LLM for multiple roles within the RAG system.
5 Conclusion

### v2

@@ 3. Experiment
3.1. Setting
3.1.1. Datasets
Following the prior works (Lin et al., 2023b; Asai et al., 2023), we choose two ODQA datasets (FreebaseQA (Yao et al., 2014) and MS-MARCO (Bajaj et al., 2016)) and one reading comprehension (RC) dataset to do the representation learning in the retrieval stage (denoted as $\mathcal{D}_{R}$) while leveraging other three datasets (Web Question (Dumais et al., 2002), Wiki Question Answering (Yang et al., 2015)) and SQuAD v2111https://rajpurkar.github.io/SQuAD-explorer/explore/v2.0/dev/Prime_number.html for fine-tuning the LLM in the generation stage (denoted as $\mathcal{D}_{G}$). The statistic of chosen datasets is shown in Tab.1. For detailed descriptions and complied templates, please refer to Appendix A.
Table 1. The Statistics of Fine-tuning Datasets.
@@ 3.1.1. Datasets
Wiki QA (Yang et al., 2015) wiki_qa 20360 Open-domain QA
FreebaseQA (Yao et al., 2014) freebase_qa 20358 Open-domain QA
MS-MARCO (Bajaj et al., 2016) ms_marco 80143 Open-domain QA
Web Question (Dumais et al., 2002) web_question 3778 Open-domain QA
@@
3.1.2. Evaluation
To access our performance, we conduct the evaluation on four knowledge-intensive datasets, such as TriviaQA (denoted as TQA) [^2] https://nlp.cs.washington.edu/triviaqa, Natural Question (denoted as NQ) [^3] https://ai.google.com/research/NaturalQuestions and PopQA [^4] https://huggingface.co/datasets/akariasai/PopQA, that are not involved in the training progress. For the evaluation metric, we evaluate our model’s generation performance using the Exact Match (Wongsuphasawat et al., 2012), which indicates whether gold answers are included in the model generations followed by the setting in prior work (Lin et al., 2023b; Mao et al., 2024). Furthermore, to evaluate our proposed retriever’s performance, we employ the Acc@5 and Acc@20 as evaluation metrics, which are widely used in related studies (Chen et al., 2024; Izacard et al., 2021). These metrics assess the proportion of questions where the correct answers appear in the top-5 or top-20 retrieval results, offering a comprehensive evaluation of the retrieval performance. For more details, please refer to the description and methods in Appdendix B.
3.1.3. Implementation Details
@@
Retrieval Stage. Following the previous work’s setting (Ma et al., 2024), as LLaMA is a decoder-only architecture, we append an end-of-sequence token <EOS> to the input sequence and regard the last layer representation as the dense representation to calculate the similarity score. Considering the possible effect caused by the size of each dense representation, we also employ the normalization procedure to map the original representation into unit vectors during both the training and inference stages. For the fine-tuning progress in the retrieval stage, we adopt LoRA architecture (Hu et al., 2021) to reduce the high cost of GPU memory. The detailed hyperparameters we used can be found in Appendix C.
Generation Stage. We hold $\overline{k}$ in Sec.2.3 equal to 5 to generate instances for a single example and append multiple examples together to improve the efficiency (the length is limited to 4096 tokens). The used hyperparameters are also shown in the Appendix C. Other implementation details are the same as original papers (Lin et al., 2023b; Shi et al., 2023).
3.1.4. Baselines
To demonstrate the effectiveness of our proposed architecture, we compare the retrieval performance of our Invar-RAG with state-of-the-art retrieval methods, including sparse retrieval (BM25 (Ram et al., 2023)), dense retrieval (BGE (Xiao et al., 2024), Contriever (Izacard et al., 2021)) and LLM-based retrieval (LLM-embedder (Zhang et al., 2023) and RepLLaMA (Ma et al., 2024)). Furthermore, for the corresponding RAG performance, we conduct extensive experiments compared to the novel retriever + generation model to show our superiority. The descriptions for each baseline are listed in the Appendix D.
3.2. Overall Performance
@@
Table 2. Retrieval performance comparison between our designed retriever in Invar-RAG and other baselines. The best results are bold, and the second-best are underlined.
@@
BM25 (Ram et al., 2023) 62.5 73.0 49.0 67.0 35.5 51.5
BM25+BGE(re-ranker) (Chen et al., 2024) 72.5 78.0 68.0 76.5 54.0 60.0
Contriever (Izacard et al., 2021) 68.0 80.5 68.0 84.0 62.0 77.5
BGE-base (Xiao et al., 2024) 69.5 80.0 77.0 86.0 72.0 83.0
LLM-embedder (Zhang et al., 2023) 67.5 77.5 75.5 86.5 70.0 79.5
RepLLaMA (Ma et al., 2024) 66.5 76.0 72.0 85.5 68.5 74.5
@@
3.2.1. Retrieval Performance
In this section, we will present and analyze the retrieval performance of our designed architecture. As illustrated in Tab.2, the sparse retriever BM25 fails to map the given text to proper representations. Although employing an additional model as the re-ranker improves the performance to some extent, the retrieval capability remains sub-optimal due to the inferiority of BM25. Besides, Novel dense retrievers, like BGE and Contriever, present comparable performance over the three datasets, suggesting their effectiveness in leveraging contrastive learning or task-specific fine-tuning. However, they still slightly lag behind our designed Invar-retrieval because of the neglect of rich semantic information (Ma et al., 2024). Current researchers have proposed several LLM-based retrievers i.e., LLM-embedder (Zhang et al., 2023) and RepLLaMA (Ma et al., 2024), which leverage the rich prior knowledge that LLM initially has. However, due to the high cost of processing the massive corpus, it is infeasible to handle all the chunks within the LLM. Moreover, the variance problem that happens in LLM also leads to relatively inferior performance. Correspondingly, we propose our LLM-based retrieval model, Invar-retrieval, as a part of our designed Invar-RAG. The results shows that our methods outperform all the sparse, dense and LLM-based retrievers, especially under the Acc@5 measurement, contributing to our designed invariance loss in reducing the variant and ineffective patterns.
3.2.2. Generation Performance
@@
Table 3. Generation performance comparison between our designed Invar-RAG and other baselines. The best results are bold, and the second-best are underlined.
@@ 3.3. Ablation Study
Table 4. The Statistics of Datasets in Ablation study.
@@
Table 5. Ablation Study on TQA.
@@ 3.4. Invariance Analysis
  • Prior works (Zhang et al., 2024) have shown that adding a suitable amount of irrelevant or relatively ineffective documents does help improve the retrieval performance.
Figure 2. Special example for illustrating the effectiveness of invariance loss.
@@
4. Related Work
Information Retrieval: Advancements in deep learning have revolutionized information retrieval systems, enhancing their personalization and accuracy in retrieving relevant documents. Early information retrieval frameworks employed sparse retrievers (Ram et al., 2023) or dense retrievers (Izacard et al., 2021; Xiao et al., 2024) to represent large corpora but struggled to capture deep semantic relationships (Karpukhin et al., 2020). LLM-based retrievers (generative retrieval) have since emerged as notable methods, leveraging the rich prior knowledge of LLMs to significantly improve performance by converting documents into parametric knowledge and generating them instead of computing similarity scores (Zhu et al., 2021). However, the frequent encoding and decoding processes in LLMs severely hinder efficiency (Zhu et al., 2021). To address the trade-off between effectiveness and efficiency, we propose invar-retrieval in our architecture, enabling the model to efficiently retrieve the most relevant documents without introducing variance.
Retrieval-augmented Language Model: Currently, retrieval-augmented language models have proven effective in answering questions by leveraging external information through the integration of novel retrievers and LLMs (Zhu et al., 2021). However, the architectural gap between retrieval and generation continues to hinder unified optimization across the entire retrieval-augmented generation system (Fan et al., 2024). To address the isolation between retrieval and generation, a novel architecture called RA-DIT was introduced (Lin et al., 2023b). By aligning retriever scoring with LSR scoring (Shi et al., 2023), it has been shown to deliver state-of-the-art performance across various tasks. However, it still employs dense retrievers like DRAGON+ (Lin et al., 2023a) in the retrieval stage, which fails to eliminate the problem at its source and introduces inefficiencies throughout the process. Correspondingly, we introduce a representation learning method and invariance loss in our Invar-RAG architecture, which partially addresses these issues and explores a novel approach to using a single LLM for multiple roles within the RAG system.
5. Conclusion
@@
Appendix A Appendix A: Datasets Information
For training and evaluating our RAG system, we select eight datasets (five for training and three for evaluation). All the datasets we used are downloaded from the Hugging Face website [^7] https://huggingface.co/datasets.
A.1. Fine-tuning Datasets
We choose five datasets to fine-tune our two-stage architecture. Specifically, in the retrieval stage, we choose two Open-domain Question Answering (ODQA) datasets, i.e., FreebaseQA (Yao et al., 2014) and MS-MARCO (Bajaj et al., 2016), to fine-tune our model for returning more relevant documents leveraging the powerful semantic understanding capability of LLM while avoiding the variance that happened in the practical situation.
  • FreebaseQA [^8] https://huggingface.co/datasets/microsoft/freebase_qa: This dataset is designed for open-domain factoid question answering using the Freebase knowledge graph. It contains 28,348 trivia-style questions and over 54,000 question-answer matches from TriviaQA and trivia website.
  • MS-MARCO [^9] https://huggingface.co/datasets/microsoft/ms_marco: The MS-MARCO (Microsoft MAchine Reading Comprehension) is a large-scale dataset for machine reading comprehension, featuring real-world queries from Bing users and over 1 million passages with query-answer pairs, supporting tasks like question answering and passage ranking.
By using both FreebaseQA and MS-MARCO to fine-tune our retriever, we improve the retrieval part of our Invar-RAG more comprehensively. FreebaseQA provides complex, linguistically diverse queries matched with structured knowledge graph data, enhancing our model’s ability to retrieve fact-based answers. MS-MARCO, with its real-world user queries and passage ranking, improves performance in understanding and ranking everyday information. Together, they enhance the retriever’s precision, relevance, and ability to handle a variety of natural language queries. To fine-tune our model for the generation task, we select three other datasets (two ODQA datasets and one reading comprehension dataset).
  • Wiki QA [^10] https://huggingface.co/datasets/microsoft/wiki_qa: This dataset, introduced by Microsoft, consists of question and sentence pairs collected from Bing search logs. It is designed for research in open-domain question answering. The dataset contains 3,047 questions linked to Wikipedia articles, with 29,258 sentences, 1,473 of which are labeled as correct answers to the questions.
  • Web Question [^11] https://huggingface.co/datasets/Stanford/web_questions: Released as part of research on semantic parsing, the WebQuestions dataset contains 5,810 question-answer pairs where the answers are entities from Freebase. The dataset is used to train and evaluate models for answering factual questions using structured data from Freebase.
  • SQuAD v2121212https://huggingface.co/datasets/rajpurkar/squad_v2: The Stanford Question Answering Dataset (SQuAD) v2 builds upon the original SQuAD dataset by adding over 50,000 unanswerable questions that are similar to answerable ones, totaling 130,319 questions. This dataset challenges models to not only answer questions but also to determine when no answer is available in the context.
A.2. Evaluation Datasets
To evaluate both the retrieval and generation performance of Invar-RAG, we import three more ODQA datasets, i.e., TriviaQA, Natural Question, and PopQA, which are all large-scale public datasets that have been widely used as benchmarks in the retrieval or generation tasks (Fan et al., 2024).
  • TriviaQA [^13] https://huggingface.co/datasets/mandarjoshi/trivia_qa: TriviaQA is a reading comprehension dataset that comprises more than 650,000 question-answer-evidence triples. It includes 95,000 question-answer pairs generated by trivia experts, with independent evidence documents provided—around six per question on average—offering distant supervision of high quality for answering the questions.
  • Natural Question [^14] https://huggingface.co/datasets/google-research-datasets/natural_questions: The NQ dataset consists of questions sourced from real users, requiring QA systems to process and understand full Wikipedia articles that may or may not contain the answer. The use of actual user queries and the need to examine entire pages make this dataset more realistic and challenging compared to earlier QA datasets.
  • PopQA [^15] https://huggingface.co/datasets/akariasai/PopQA: PopQA is a large-scale, open-domain QA dataset with 14,000 entity-centric question-answer pairs. The questions are generated by applying templates to knowledge tuples retrieved from Wikidata.
Appendix B Appendix B: Evaluation
In this section, we will detail the information of our evaluation metrics and the implementation of our evaluation process. For retrieval, to provide a comprehensive view of performance, we leverage Accuracy both truncated at five and twenty, to assess the proportion of questions where the correct answers appear in the top 5 or top 20 retrieval results, respectively. For generations, we choose the Exact Match (Wongsuphasawat et al., 2012) to evaluate the difference between our prediction and the ground truth. Moreover, in our evaluation process, we adopt 8-shot examples which are randomly selected from the chosen datasets following previous work (Izacard & Grave, 2020).
Appendix C Appendix C: Hyperparameter Illustration
In this section, we will present the prime hyperparameter we used and the values we set in our fine-tuning and inferencing progress. Tab.6 shows the hyperparameter for fine-tuning our retrieval stage and generation stage.
Table 6. Hyperparameter for Fine-tuning (Retrieval and Generation)
Stage lr_scheduler batch size seq len LoRA_rank LoRA_alpha LoRA_dropout
Retrieval Stage cosine 64 4096 16 32 0.05
Generation Stage cosine 64 4096 16 32 0.05
Appendix D Appendix D: Baselines Description
(a) BM25 (Ram et al., 2023): A classical, sparse information retrieval method based on term frequency and inverse document frequency (TF-IDF). It ranks documents based on the occurrence of query terms, emphasizing exact term matches, making it effective in domains with well-structured data but limited in understanding semantic similarities. (b) BGE (Izacard et al., 2021): A family of pre-trained embedding models designed for general Chinese text embedding. It achieves superior performance across diverse tasks by using massive datasets and advanced training techniques like contrastive learning, supporting retrieval, ranking, and classification. (c) Contriever (Xiao et al., 2024): Contriever is an unsupervised dense information retriever based on contrastive learning. It is designed to overcome limitations in traditional sparse retrieval methods like BM25, particularly in zero-shot and multilingual settings. Contriever demonstrates state-of-the-art performance on multiple retrieval benchmarks, such as BEIR, and excels in cross-lingual retrieval. Its architecture relies on a bi-encoder model, which encodes queries and documents independently, and uses unsupervised contrastive learning to optimize retrieval, even without labeled data. (d) LLM-embedder (Zhang et al., 2023): A unified embedding model optimized for retrieval augmentation in large language models (LLMs). It leverages multi-task fine-tuning and LLM feedback to retrieve relevant knowledge, tools, and memory, enhancing LLMs’ capabilities in knowledge-intensive tasks. (e) RepLLaMA (Ma et al., 2024): RepLLaMA is a dense retriever fine-tuned from the LLaMA-2 language model, part of a multi-stage retrieval pipeline that includes rankLLaMA as a reranker. RepLLaMA improves the effectiveness of retrieval tasks, particularly in passage and document retrieval. It outperforms smaller models, showcasing strong generalization and effectiveness in both in-domain and zero-shot settings. The model leverages both hard negatives and in-batch negatives during training.
Appendix E Appendix E: Ablation Study
In this section, we will present the ablation study by gradually removing the significant components of our architecture. Specifically, we follow the variant settings in the main body and perform our ablation study on the other two evaluation ODQA datasets (Natural Question  [^16] https://huggingface.co/datasets/google-research-datasets/natural_questions and PopQA  [^17] https://huggingface.co/datasets/akariasai/PopQA), presented in Tab.7.
Table 7. Ablation Study on Natural Question and PopQA.
Model Variants Nature Question PopQA
Retrieval Generation Retrieval Generation
Acc@5 Acc@20 Exact Match Acc@5 Acc@20 Exact Match
Default 80.5 88.0 56.2 73.5 82.5 53.6
w/o representation learning 69.0 80.5 54.7 59.5 70.0 51.7
w/o invariance loss 77.5 86.0 55.4 69.5 81.0 52.4
w/o generative fine-tuning / / 53.9 / / 50.7
From the experimental results, we can see that the performance of our Invar-RAG, no matter for the retrieval and the generation, shows a similar tendency to the one on TriviaQA, showing the effectiveness of our designed two-stage fine-tuning method, featuring the representation learning, invariance loss and the generation fine-tuning in one single LLM.